### v2

@@ 4 Mitigating Implicit Penalty by Residual Learning: Tiki-Taka
Under the strongly convex assumption, the optimal point $W^{*}$ is unique. We believe the requirement of strong convexity is non-essence, and the proof can be extended to more general cases, which is left for future work.
Involving two sequences $W_{k}$ and $P_{k}$, Tiki-Taka converges in different senses, including: (a) the residual array $P_{k}$ converges to the optimal point $P^{*}(W_{k})$ of (15); (b) $W_{k}$ converges to the critical point of (18) or the optimal point $W^{*}$; (c) the sum $\bar{W}_{k}=W_{k}+\gamma P_{k}$ converges to a critical point where $\nabla f(\bar{W}_{k})$. Taking all these into account, we define the convergence metric as
$E^{\texttt{TT}}_{K}:=\frac{1}{K}\sum_{k=0}^{K-1}$ $\ {\mathbb{E}}\bigg{[}\|\nabla f(\bar{W}_{k})\|^{2}+O(\left\|P_{k}-P^{*}(W_{k})\right\|^{2})$
$\ +O(\|W_{k}-W^{*}\|^{2})\bigg{]}.$ (19)
For simplicity, the constants in front of some terms in $E^{\texttt{TT}}_{K}$ are hidden. Now, we provide the convergence of Tiki-Taka with generic responses.
CIFAR10 CIFAR100
DSGD ASGD TT DSGD ASGD TT
ResNet18 95.43 $\pm$0.13 84.47 $\pm$3.40 94.81 $\pm$0.09 81.12 $\pm$0.25 68.98 $\pm$1.01 76.17 $\pm$0.23
ResNet34 96.48 $\pm$0.02 95.43 $\pm$0.12 96.29 $\pm$0.12 83.86 $\pm$0.12 78.98 $\pm$0.55 80.58 $\pm$0.11
ResNet50 96.57 $\pm$0.10 94.36 $\pm$1.16 96.34 $\pm$0.04 83.98 $\pm$0.11 79.88 $\pm$1.26 80.80 $\pm$0.22
Table 1: Analog training with the power response for fine-tuning task on CIFAR10/100 using models from ResNet family. The test accuracy is reported. DSGD, ASGD, and TT represent Digital SGD, Analog SGD, Tiki-Taka, respectively.
Theorem 4.2 (Convergence of Tiki-Taka).
Under Assumptions 3.23.4, and 4.1, with learning rate $\alpha=O\left(\sqrt{{1}/{\sigma^{2}K}}\right)$, $\beta=O(\alpha\gamma^{3/2})$, it holds for Tiki-Taka that
$E^{\texttt{TT}}_{K}\leq\ O\left(\sqrt{{\sigma^{2}}/{K}}+\sigma^{2}S^{\texttt{TT}}_{K}\right)$ (20)
where $S^{\texttt{TT}}_{K}$ denotes the amplification factor of $P_{k}$ given by $S^{\texttt{TT}}_{K}:=\frac{1}{K}\sum_{k=0}^{K}\left\|\frac{G(P_{k})}{\sqrt{F(P_{k})}}\right\|^{2}_{\infty}$.
The proof of Theorem 4.2 is deferred to Appendix G. Theorem 4.2 claims that Tiki-Taka converges at the rate $O\left(\sqrt{{\sigma^{2}}/{K}}\right)$ to a neighbor of critical point with radius $O(\sigma^{2}S^{\texttt{TT}}_{K})$, which share almost the same expression with the convergence of Analog SGD. The difference lies in the amplification factor $S^{\texttt{TT}}_{K}$ and $S^{\texttt{ASGD}}_{K}$, where the former depends on $P_{k}$ while the latter depends on $W_{k}$.
Impact of response functions. Response function affects the Analog SGD and Tiki-Taka similarly. However, attributed to the residual array, constructing response functions to enable exact convergence of Tiki-Taka is feasible.
As we have discussed, $P_{k}$ tends to $P^{*}(W_{k})$ which tends to 00 given $W_{k}$ tends to $W^{*}$. Resistive elements with response functions such that $G(P)=0$ when $P=0$ is required for the exact convergence.
Assumption 4.3.
@@
Corollary 4.4 (Exact convergence of Tiki-Taka).
Under Assumption 4.3 and the conditions in Theorem 4.2, if $\gamma\geq\Omega(H_{\min}^{-1/5})$, it holds that $E^{\texttt{TT}}_{K}\leq\ O\left(\sqrt{{\sigma^{2}L}/{K}}\right)$.
The proof of Corollary 4.4 is deferred to Appendix G.5. Corollary 4.4 demonstrates the failure of Tiki-Taka in Figure 2. The symmetric point is $w^{\diamond}=c_{\texttt{Lin}}\tau$ in this example, which violates Assumption 4.3 when $c_{\texttt{Lin}}\neq 0$ and hence introduces asymptotic error into Tiki-Taka.
@@ 5 Numerical Simulations
In this section, we verify the main theoretical results by simulations on both synthetic datasets and real datasets. We use the open source toolkit AIHWKIT (AIHWKIT(Rasch et al., 2021) to simulate the behaviors of Analog SGD and Tiki-Taka. Each simulation is repeated three times, and the mean and standard deviation are reported. More details can be referred to in Appendix I.
We consider two types of response functions in our simulations: power and exponential response functions with dynamic ranges $[-\tau,\tau]$, while the symmetric point is 0, as required by Corollary 4.4.
MNIST FCN/CNN. We train fully-connected network (FCN) and convolution neural network (CNN) on the MNIST dataset and compare the performance of Analog SGD and Tiki-Taka under various $\tau$ on power responses; see the results in Figure 4. By tracking residual, Tiki-Taka outperforms Analog SGD and reaches comparable accuracy with Digital SGD. For both architectures, the accuracy of Tiki-Taka drops by $<1\%$. In contrast, Analog SGD takes a few epochs to achieve observable accuracy increment in FCN training, rendering a slower convergence rate than Tiki-Taka. In CNN training, Analog SGD’s accuracy increases more slowly than Tiki-Taka does and finally gets stuck at about 80%. It is consistent with the theoretical claims.
CIFAR10/CIFAR100 ResNet. We fine-tune three models from the ResNet family with different scales on CIFAR10/CIFAR100 datasets. The power response functions with the parameter $\gamma_{\text{res}}=3.0$ and $\tau=0.1$ are used, whose results are shown in Table 1. The results show that the Tiki-Taka outperforms Analog SGD by about $1.0\%$ in most of the cases in ResNet34/50, and the gap even reaches about $10.0\%$ for ResNet18 training on the CIFAR100 dataset.
@@ 6 Conclusions and Limitations
This paper studies the impact of response functions on gradient-based training in analog in-memory computing hardware. We first formulate the dynamic of Analog Update based on the pulse update rule, which is a bridge between the equation at the resistive element level and the one at the algorithm level. Based on this dynamic, we study the convergence of two gradient-based analog training algorithms, Analog SGD and Tiki-Taka. The theoretical results demonstrate that Analog SGD converges to a critical point inexactly with asymptotic error, which comes from the noise and asymmetric update. To overcome this issue, Tiki-Taka introduces a residual array whose stationary point is 0. By properly shifting the symmetric point of the residual array, Tiki-Taka provably converges to a critical point exactly. Numerical simulations verify the claims about the implicit bias and the efficacy of Tiki-Taka against Analog SGD.
@@ Appendix A Literature Review
Resistive element. A series of works seek various resistive elements that have near-constant or at least symmetric responses. The leading candidates currently include PCM (Burr et al., 2016; Le Gallo & Sebastian, 2020), ReRAM (Jang et al., 2014; 2015; Stecconi et al., 2024), CBRAM (Lim et al., 2018; Fuller et al., 2019), ECRAM (Tang et al., 2018; Onen et al., 2022), MRAM (Jung et al., 2022; Xiao et al., 2024), FTJ (Guo et al., 2020) or flash memory (Wang et al., 2018a; Xiang et al., 2020; Merrikh-Bayat et al., 2017).
@@
However, the resistive element possessing symmetric updates may not be the best option in terms of manufacturing. For example, although ECRAM provides almost symmetric updates, it is still less competitive than ReRAM as ReRAM has a faster response speed and lower pulse voltage (Stecconi et al., 2024). The suitability of the resistive elements is evaluated based on metrics across multiple dimensions, such as the number of conductance states, retention, material endurance, switching energy, response speed, manufacturing cost, and cell size. Among them, this paper is only interested in the impact of response functions in the training.
Gradient-based training on AIMC hardware. A series of works focuses on implementing back-propagation (BP) and gradient-based training on AIMC hardware. The seminal work (Gokmen & Vlasov, 2016; Gokmen et al., 2017) leverages the rank-one structure of the gradient and implements Analog SGD by a stochastic pulse update scheme, rank-update. Rank-update significantly accelerates the gradient descent step by avoiding dealing the gradients with $O(N^{2})$ elements directly but using two $O(N)$ vectors for update instead, where $N$ is the numbers of matrix row and column. To alleviate the asymmetric update issue, researchers also design various of Analog SGD variants, Tiki-Taka algorithm family (Gokmen & Haensch, 2020; Gokmen, 2021; Rasch et al., 2024). The key components of Tiki-Taka are introducing a residual array to stabilize the training. Apart from the rank-update, a hybrid scheme that performs forward and backward in the analog domain but computes the gradients in the digital domain has been proposed in (Nandakumar et al., 2018; 2020). Their solution, referred to as mixed-precision update, provides a more accurate gradient signal but requires 5$\times$-10$\times$ higher overhead compared to the rank-update scheme (Rasch et al., 2024).
@@
Attributed to these efforts, analog training has empirically shown great promise in achieving a similar level of accuracy as digital training on chip prototype, with reduced energy consumption and training time (Wang et al., 2018b; Gong et al., 2022). Simultaneously, the parallel acceleration solution with AIMC hardware is also under exploration (Wu et al., 2024b). Despite its good performance, it is still mysterious about when and why they work.
@@
Theoretical foundation of gradient-based training. The closely related result comes from the convergence study of Tiki-Taka (Wu et al., 2024a). Similar to our work, they attempt to model the dynamic and provide the convergence properties of Analog SGD and Tiki-Taka. However, their work is limited to a special linear response function. Furthermore, their paper considers a simplified version of Tiki-Taka, with a hyper-parameter $\gamma=0$ (see Section 4). As we will show empirically and theoretically, Tiki-Taka benefits from a non-zero $\gamma$. Consequently, We compare the results briefly in Table 2 and comprehensively in Appendix B.
@@
Tiki-Taka (Wu et al., 2024a) $=0$ $O\left(\sqrt{\frac{1}{K}}\frac{1}{1-33P_{\max}^{2}/\tau^{2}}\right)$
Tiki-Taka  [Corollary 4.4] $\neq 0$ $O\left(\sqrt{\frac{1}{K}}\frac{1}{H^{\texttt{TT}}_{\min}}\right)$ $O\left(\sqrt{\frac{1}{K}}\frac{1}{1-P_{\max}^{2}/\tau^{2}}\right)$
Table 2: Comparison between our paper and Wu et al. (2024a). Mixing-coefficient $\gamma$ is a hyper-parameter of Tiki-Taka. “Generic response” and “Linear response” columns are the convergence rates in the corresponding settings. $K$ represents the number of iterations. $H^{\texttt{TT}}_{\min}$ and $P_{\max}^{2}/\tau^{2}<1$ measure the saturation while the former one reduces to the latter on linear response functions.
Energy-based models and equilibrium propagation. Apart from achieving explicit gradient signals by the BP, there are also attempts to train models based on equilibrium propagation (EP, Scellier & Bengio (2017)), which provides a biologically plausible alternative to traditional BP. EP is applicable to a series of energy-based models, where the forward pass is performed by minimizing an energy function (Watfa et al., 2023; Scellier et al., 2024). The update signal in EP is computed by measuring the output difference between a free phase and an active phase. EP eliminates the need for BP non-local weight transport mechanism, making it more compatible with neuromorphic and energy-efficient hardware (Kendall et al., 2020; Ernoult et al., 2020). We highlight here that the approach to attain update signals (BP or EP) is orthogonal to the update mechanism (pulse update). Their difference lies in the objective $f(W_{k})$, which is hidden in this paper. Therefore, building upon the pulse update, our work is applicable to both BP and EP.
Physical neural network. The model executing on AIMC hardware, which leverages resistive crossbar array to accelerate MVM operation, is a concrete implementation of physical neural networks (PNNs, Wright et al. (2022); Momeni et al. (2024)). PNN is a generic concept of implementing neural networks via a physical system in which a set of tunable parameters, such as holographic grating (Psaltis et al., 1990), wave-based systems (Hughes et al., 2019), and photonic networks (Tait et al., 2017). Our work particularly focuses on training with AIMC hardware, but the methodology developed in this paper can be transferred to the study of other PNNs.
Appendix B Relation with the result in Wu et al. (2024a)
Similar to this paper, Wu et al. (2024a) also attempts to model the dynamic of analog training. They shows that Analog SGD converges to a critical point of problem (1) inexactly with an asymptotic error, and Tiki-Taka converges to a critical point exactly. In this section, we compare our results with our results and theirs.
As discussed in Section 1, Wu et al. (2024a) studies the analog training on special linear response functions

### v3

@@ 4 Mitigating Implicit Penalty by Residual Learning
Under the strongly convex assumption, the optimal point $W^{*}$ is unique. Although the requirement of strong convexity is non-essential by the development of bilevel optimization [40, 41, 42, 43], and the proof can be extended to more general cases, we leave it for future work.
Involving two sequences $W_{k}$ and $P_{k}$, Residual Learning converges in different senses, including: (a) the residual array $P_{k}$ converges to the optimal point $P^{*}(W_{k})$; (b) $W_{k}$ converges to the critical point of $f(\cdot)$ or the optimal point $W^{*}$; (c) the sum $\bar{W}_{k}=W_{k}+\gamma P_{k}$ converges to a critical point where $\nabla f(\bar{W}_{k})=0$. Taking all these into account, we define the convergence metric as
$E^{\texttt{RL}}_{K}:=\frac{1}{K}\sum_{k=0}^{K-1}$ $\ {\mathbb{E}}\bigg[\|\nabla f(\bar{W}_{k})\|^{2}+O(\left\|P_{k}-P^{*}(W_{k})\right\|^{2})+O(\|W_{k}-W^{*}\|^{2})\bigg].$ (12)
For simplicity, the constants in front of some terms in $E^{\texttt{RL}}_{K}$ are hidden. Now, we provide the convergence of Residual Learning with generic responses.
Theorem 3 (Convergence of Residual Learning).
Under Assumptions 13 and 4, with the learning rate $\alpha=O\left(\sqrt{{1}/{\sigma^{2}K}}\right)$, $\beta=O(\alpha\gamma^{3/2})$, it holds for Residual Learning that
$E^{\texttt{RL}}_{K}\leq\ O\left(\sqrt{{\sigma^{2}}/{K}}+\sigma^{2}S^{\texttt{RL}}_{K}\right)$ (13)
where $S^{\texttt{RL}}_{K}$ denotes the amplification factor of $P_{k}$ given by $S^{\texttt{RL}}_{K}:=\frac{1}{K}\sum_{k=0}^{K}\left\|\frac{G(P_{k})}{\sqrt{F(P_{k})}}\right\|^{2}_{\infty}$.
The proof of Theorem 3 is deferred to Appendix I. Theorem 3 claims that Residual Learning converges at the rate $O\left(\sqrt{{\sigma^{2}}/{K}}\right)$ to a neighbor of critical point with radius $O(\sigma^{2}S^{\texttt{RL}}_{K})$, which share almost the same expression with the convergence of Analog SGD. The difference lies in the amplification factor $S^{\texttt{RL}}_{K}$ and $S^{\texttt{ASGD}}_{K}$, where the former depends on $P_{k}$ while the latter depends on $W_{k}$.
Impact of response functions. Response function affects the Analog SGD and Residual Learning similarly. However, attributed to the residual array, constructing response functions to enable exact convergence of Residual Learning is viable.
As we have discussed, $P_{k}$ tends to $P^{*}(W_{k})$ which tends to 0 given $W_{k}$ tends to $W^{*}$. Therefore, response functions with $G(P)=0$ when $P=0$ are required for the exact convergence.
Assumption 5.
@@
Corollary 1 (Exact convergence of Residual Learning).
Under Assumption 5 and the conditions in Theorem 3, if $\gamma\geq\Omega(R_{\min}^{-1/5})$, it holds that $E^{\texttt{RL}}_{K}\leq\ O\left(\sqrt{{\sigma^{2}L}/{K}}\right)$.
The proof of Corollary 1 is deferred to Appendix I.5. Corollary 1 demonstrates the failure of Tiki-Taka in Figure 2. The symmetric point is $w^{\diamond}=c_{\texttt{Lin}}\tau$ in this example, which violates Assumption 5 when $c_{\texttt{Lin}}\neq 0$ and hence introduces asymptotic error into Residual Learning.
Figure 4: Test accuracy of training on MNIST dataset under different $\tau$; (Left) FCN. (Right) CNN.
5 Extension of Residual Learning: limited granularity and noisy IO
This section extends Residual Learning to more practical scenarios involving more hardware imperfections. To be specific, we consider the noisy IO and limited granularity as examples. We highlight that we are not trying to diminish the importance of imperfection, but we focus on two of the primary ones that are known to be crucial.
Noisy IO introduces noise during the reading of $P_{k+1}$ in the transfer process (11), given by $W_{k+1}=W_{k}+\beta(P_{k+1}+\varepsilon_{k})\odot F(W_{k})-\beta|P_{k+1}+\varepsilon_{k}|\odot G(W_{k})$ with a noise $\varepsilon_{k}$. It incurs the implicit penalty issues again, leading to a penalized upper-level objective $\|P^{*}(W)\|^{2}+\Sigma_{\varepsilon}\|W\|^{2}/2$, as claimed by Theorem˜1, where $\Sigma_{\varepsilon}$ is the standard variation of $\varepsilon_{k}$. To filter out the noise, we propose to use a digital buffer $H_{k}$ to take a moving average of noisy $P_{k+1}$ signals by
$H_{k+1}=$ $\ (1-\beta)H_{k}+\beta(P_{k+1}+\varepsilon_{k+1})$ (14)
Intuitively, with a fixed $P_{k+1}$, $H_{k}$ will converge to a neighborhood of $P_{k+1}$ with radius $O(\beta)$. Therefore, a sufficiently small $\beta$ renders $H_{k}$ a fair approximation of noiseless $P_{k}$, enabling optimizing the upper-level objective with clearer signals. After that, $H_{k+1}$ is transferred to $W_{k}$ as follows
$W_{k+1}=$ $\ W_{k}+\beta H_{k+1}\odot F(W_{k})-\beta|H_{k+1}|\odot G(W_{k}).$ (15)
Furthermore, the optimization process suffers from a constant error in the order of $O(\Delta w_{\min})$ as the pulses are fired discretely, with each changing the weight by $O(\Delta w_{\min})$. To overcome these issues, we propose introducing a threshold mechanism to train in limited precision. Instead of transferring $H_{k+1}$ to $W_{k}$ at each iteration, the transfer is triggered only if an element in $H_{k+1}$ is greater than the response granularity $\Delta w_{\min}$. The proposed algorithms are referred to as Residual Learning v2.
CIFAR10
DSGD ASGD TT/RL TTv2 RLv2
ResNet18 95.43 $\pm$0.13 84.47 $\pm$3.40 94.81 $\pm$0.09 95.31 $\pm$0.05 95.12 $\pm$0.14
ResNet34 96.48 $\pm$0.02 95.43 $\pm$0.12 96.29 $\pm$0.12 96.60 $\pm$0.05 96.42 $\pm$0.13
ResNet50 96.57 $\pm$0.10 94.36 $\pm$1.16 96.34 $\pm$0.04 96.63 $\pm$0.09 96.56 $\pm$0.08
CIFAR100
DSGD ASGD TT/RL TTv2 RLv2
ResNet18 81.12 $\pm$0.25 68.98 $\pm$1.01 76.17 $\pm$0.23 78.56 $\pm$0.29 79.83 $\pm$0.13
ResNet34 83.86 $\pm$0.12 78.98 $\pm$0.55 80.58 $\pm$0.11 81.81 $\pm$0.15 82.85 $\pm$0.19
ResNet50 83.98 $\pm$0.11 79.88 $\pm$1.26 80.80 $\pm$0.22 82.82 $\pm$0.33 83.90 $\pm$0.20
Table 1: Fine-tuning ResNet models with the power response on CIFAR10/100. Test accuracy is reported. DSGD, ASGD, TT/RL, TTv2, and RLv2 represent Digital SGD, Analog SGD, Residual Learning/Tiki-Taka, and Residual Learning v2, respectively.
@@ 6 Numerical Simulations
In this section, we verify the main theoretical results by simulations on both synthetic datasets and real datasets. We use the open source toolkit AIHWKIT (AIHWKIT[44] to simulate the behaviors of Analog SGD, Residual Learning (which reduces to Tiki-Taka). Each simulation is repeated three times, and the mean and standard deviation are reported. We consider two types of response functions in our simulations: power and exponential response functions with dynamic ranges $\tau,\tau]$, while the symmetric point is 0, as required by Corollary 1. More details, simulations, and ablation studies can be found in Appendix K.
MNIST FCN/CNN. We train a fully-connected network (FCN) and a convolutional neural network (CNN) on the MNIST dataset and compare the performance of Analog SGD and Tiki-Taka under various $\tau$ on power responses; see the results in Figure 4. By tracking residual, Tiki-Taka outperforms Analog SGD and reaches comparable accuracy with Digital SGD. For both architectures, the accuracy of Tiki-Taka drops by $1\%$. In contrast, Analog SGD takes a few epochs to achieve an observable accuracy increment in FCN training, rendering a slower convergence rate than Tiki-Taka. In CNN training, Analog SGD’s accuracy increases more slowly than Tiki-Taka does and finally gets stuck at about 80%. It is consistent with the theoretical claims.
CIFAR10/CIFAR100 ResNet. We fine-tune three ResNet models with different scales on CIFAR10/CIFAR100 datasets. The power response functions are used, whose results are shown in Table 1. The results show that the Tiki-Taka outperforms Analog SGD by about $1.0\%$ in most of the cases in ResNet34/50, and the gap even reaches about $7.0\%$ for ResNet18 training on the CIFAR100 dataset. On top of that, we also compare the proposed Residual Learning v2 and Tiki-Taka v2. Both of them outperform Residual Learning since they introduce a digital buffer to filter out the reading noise. However, Residual Learning v2 outperforms Tiki-Taka v2 on the CIFAR100 dataset, demonstrating the benefit from the bilevel formulation.
@@ 7 Conclusions and Limitations
This paper studies the impact of a generic class of asymmetric and non-linear response functions on gradient-based training in analog in-memory computing hardware. We first formulate the dynamic of Analog Update based on the pulse update rule. Based on it, we show that Analog SGD implicitly optimizes a penalized objective and hence can only converge inexactly. To overcome this issue, we propose a Residual Learning framework which solves a bilevel optimization problem. Explicitly aligning the algorithmic stationary point and physical symmetric point, Residual Learning provably converges to the optimal point exactly. Furthermore, we demonstrate how to extend Residual Learning to overcome the noisy reading and limited update granularity issues. The efficiency of the proposed method is verified through simulations. One limitation of this work is that the current analysis considers only the three hardware imperfections. While they are known to be crucial for analog training, it is also important to extend our convergence analysis and methods to more practical scenarios involving more imperfections in future work.
@@ Appendix A Literature Review
Training on AIMC hardware. Analog training has shown promising early successes in certain tasks like face classification [25] and digit classification [26], where $1,000\times$ lower energy consumption compared with digital implementation is presented. Researchers are also exploring approaches to mitigate the impact of hardware non-idealities. For example, [27, 28] proposes to leverage the momentum technique to stabilize the training by reducing the noise. In order to overcome other potential non-idealities, a hybrid training paradigm is also being explored. [29] leverages the chip-in-the-loop technique to train models layer-by-layer, while [30] proposes to train the backbone in the digital domain and train the last layer in the analog domain. In general, these works have provided valuable insights into analog training, shedding light on many critical technical challenges. However, their focus has largely been on experimental and simulation aspects, with limited systematic and theoretical analysis of how specific imperfections affect the training process. In our paper, we present an alternative viewpoint and novel tools to explore the effects of non-idealities.
@@
Resistive element. A series of works seeks various resistive elements that have near-constant or at least symmetric responses. The leading candidates currently include PCM [45, 46], ReRAM [47, 48, 49], CBRAM [50, 51], ECRAM [52, 53], MRAM [54, 55], FTJ [56] or flash memory [57, 58, 59].
@@
However, the resistive element possessing symmetric updates may not be the best option in terms of manufacturing. For example, although ECRAM provides almost symmetric updates, it is still less competitive than ReRAM as ReRAM has a faster response speed and lower pulse voltage [49]. The suitability of the resistive elements is evaluated based on metrics across multiple dimensions, such as the number of conductance states, retention, material endurance, switching energy, response speed, manufacturing cost, and cell size. Among them, this paper is only interested in the impact of response functions in the training.
Imperfection of AIMC hardware. Besides the response functions, analog training suffers from all kinds of hardware imperfection, especially when the task’s scale increases, like asymmetric update [17, 19], reading/writing noise [18, 60, 61], device/cycle variations [62], non-linear current response due to IR drop [18, 6, 63]. This paper mainly focuses on asymmetric response functions. However, this paper is not trying to diminish the importance of other hardware imperfections but focuses on one of the primary ones that are known to be very important [19, 16].
Hardware-aware training. For inference on AIMC hardware purposes, models pretrained on digital hardware will be programmed on analog hardware. Due to the hardware imperfection, the pretrained models suffer from performance drops. Hardware-aware training (HWA) is a technique designed to bridge the gap between ideal pretrained models and non-ideal programmed models. In contrast to standard training methods, hardware-aware training explicitly incorporates device-specific imperfections, such as weight drift [20], device fail [64], bounded dynamic range [65], quantization error from ADC [66, 67, 68], device variation [69], and non-linear current output [70], into the training loop. By modeling these constraints during the training phase, the learned parameters become inherently more robust to real-world deployment conditions. It is worth highlighting that HWA is still performed on digital hardware, and the trained model will be programmed onto AIMC hardware. On the contrary, this paper considers a different and more challenging setting where the training is performed directly on analog hardware.
[…]
@@
Attributed to these efforts, analog training has empirically shown great promise in achieving a similar level of accuracy as digital training on chip prototype, with reduced energy consumption and training time [72, 73]. Simultaneously, the parallel acceleration solution with AIMC hardware is also under exploration [74]. Despite its good performance, it is still mysterious about when and why they work.
@@
Theoretical foundation of gradient-based training. The closely related result comes from the convergence study of Tiki-Taka [21]. Similar to our work, they attempt to model the dynamic and provide the convergence properties of Analog SGD and Tiki-Taka. However, their work is limited to a special linear response function. Furthermore, their paper considers a simplified version of Tiki-Taka, with a hyper-parameter $\gamma=0$ (see Section˜4). As we will show empirically and theoretically, Tiki-Taka benefits from a non-zero $\gamma$. Consequently, We compare the results briefly in Table 2 and comprehensively in Appendix B.
@@
Tiki-Taka [21] =0=0 $O\left(\sqrt{\frac{1}{K}}\frac{1}{1-33P_{\max}^{2}/\tau^{2}}\right)$
Tiki-Taka  [Corollary 1] $\neq 0$ $O\left(\sqrt{\frac{1}{K}}\frac{1}{R^{\texttt{RL}}_{\min}}\right)$ $O\left(\sqrt{\frac{1}{K}}\frac{1}{1-P_{\max}^{2}/\tau^{2}}\right)$
Table 2: Comparison between our paper and [21]. Mixing-coefficient $\gamma$ is a hyper-parameter of Tiki-Taka. “Generic response” and “Linear response” columns are the convergence rates in the corresponding settings. $K$ represents the number of iterations. $R^{\texttt{RL}}_{\min}$ and $P_{\max}^{2}/\tau^{2}<1$ measure the saturation while the former one reduces to the latter on linear response functions.
Energy-based models and equilibrium propagation. Apart from achieving explicit gradient signals by the BP, there are also attempts to train models based on equilibrium propagation (EP, [33]), which provides a biologically plausible alternative to traditional BP. EP is applicable to a series of energy-based models, where the forward pass is performed by minimizing an energy function [34, 35]. The update signal in EP is computed by measuring the output difference between a free phase and an active phase. EP eliminates the need for BP non-local weight transport mechanism, making it more compatible with neuromorphic and energy-efficient hardware [36, 37]. We highlight here that the approach to attain update signals (BP or EP) is orthogonal to the update mechanism (pulse update). Their difference lies in the objective $f(W_{k})$, which is hidden in this paper. Therefore, building upon the pulse update, our work is applicable to both BP and EP.
Physical neural network. The model executing on AIMC hardware, which leverages resistive crossbar array to accelerate MVM operation, is a concrete implementation of physical neural networks (PNNs, [75, 76]). PNN is a generic concept of implementing neural networks via a physical system in which a set of tunable parameters, such as holographic grating [77], wave-based systems [78], and photonic networks [79]. Our work particularly focuses on training with AIMC hardware, but the methodology developed in this paper can be transferred to the study of other PNNs.
Appendix B Relation with the result in [21]
Similar to this paper, [21] also attempts to model the dynamic of analog training. They shows that Analog SGD converges to a critical point of problem (1) inexactly with an asymptotic error, and Tiki-Taka converges to a critical point exactly. In this section, we compare our results with our results and theirs.
As discussed in Section 1, [21] studies the analog training on special linear response functions